Demography & International Capital Flows^{∗}

David Backus, Thomas Cooley, and Espen Henriksen Revised: May 27, 2011

This is very much research in progress. There are two related questions to which we don’t yet have answers, but we think we have an approach, some suggestive data, and an idea of what the problems are.

1. Why are savings rates in China so high? Can they be explained by demography and standard life-cycle models given current policies? Answer. We can certainly explain some of it. Demography does a good job of explaining variations in savings rates in G8 countries, including the US and Japan. We are still working on China, whose saving rates are extraordinarily high.

2. Why is China exporting capital when it seems to need capital internally? We have only clues BUT one can only address this question in a general equilibrium context in which you take account of demographics in all the relevant countries. What matters for capital flows and net foreign asset positions are relative demographics. But other things also matter: total factor productivity. Other wedges of inefficiency that affect investment decisions in a give country and capital labor ratios etc.

At this stage we can deliver a clear picture of demographics and capital positions. We can account for a large part of the changes in asset positions with demographics. China is a work in progress.

∗Notes for a presentation at David Denoon’s NYU Conference on Chinese Capital Markets, May 31, 2011.

The content reflects a current project that explores the sources of international capital flows and the role demography in particular. We thank the NYU Center on US-China Relations for its support.